OpenStudy (anonymous):

airplane flying in a circular path with a radius of 2 miles at a rotational speed of 3 radians per minute. find the speed (in mph)?

OpenStudy (phi):

you need to find out a few numbers: 1) how many radians is a full circle ? i.e. how many radians to equal 360º

OpenStudy (anonymous):

there are 2pi radians in the circle so 2pi=2 x 3.14 = 6.28

OpenStudy (phi):

yes 2 pi radians in a circle so we can figure out how many minutes it takes the plane to go 2 pi radians: \[ \frac{2 \pi}{3} \text{ minutes}\] next we need to know how far the plane traveled when it went around the circle once. \[ C = 2\pi r = 2 \cdot 2 \pi \text{ miles} = 4 \pi \text{ miles}\]

OpenStudy (phi):

distance = speed * time and speed = distance / time that means we can find the speed in miles per minute \[ speed = \frac{ 4 \pi}{\frac{2 \pi}{3}} = 4\pi \frac{3}{2\pi}= 6 \frac{ miles}{minute}\]

OpenStudy (phi):

to change that to miles per hour multiply by 60 minutes per hour \[ 6 \frac{miles}{\cancel{minute}} \cdot \frac{60 \cancel{minute}}{1 \text{ hour}} \]

OpenStudy (anonymous):

so that will be 360 mph

OpenStudy (anonymous):

Thank you so much for the detailed answer, I really appreciate your effort. thank you @phi